1.112.1 Fundamentals of TCP/IP Weight 4

Linux Professional Institute Certification — 102

Geoffrey Robertson ge@ffrey.com

               Nick Urbanik nicku@nicku.org

2005 July

Description of Objective

Candidates should demonstrate a proper understanding of network fundamentals. This objective includes the understanding of IP-addresses, network masks and what they mean (i.e. determine a network and broadcast address for a host based on its subnet mask in "dotted quad" or abbreviated notation or determine the network address, broadcast address and netmask when given an IP-address and number of bits). It also covers the understanding of the network classes and classless subnets (CIDR) and the reserved addresses for private network use. It includes the understanding of the function and application of a default route. It also includes the understanding of basic internet protocols (IP, ICMP, TCP, UDP) and the more common TCP and UDP ports (20, 21, 23, 25, 53, 80, 110, 119, 139, 143, 161).

Key files, terms, and utilities include:

/etc/services

\underline{2}

ftp
telnet
host
ping
dig
traceroute
whois

(2.2) Networking Fundamentals [14]

1.112.1

Fundamentals of TCP/IP [4]

1.112.3 TCP/IP configuration and troubleshooting [7]

1.112.4 Configure Linux as a PPP client [3]

Fundamentals of TCP/IP [4] Candidates should demonstrate a proper understanding of network fundamentals. This objective includes the understanding of IP-addresses, network masks and what they

mean (i.e. determine a network and broadcast address for a host based on its subnet mask in "dotted quad" or abbreviated notation or determine the network address, broadcast address and netmask when given an IP-address and number of bits).

It also covers the understanding of the network classes and classless subnets (CIDR) and the reserved addresses for private network use.

It includes the understanding of the function and application of a default route.

It also includes the understanding of basic internet protocols (IP, ICMP, TCP, UDP) and the more common TCP and UDP ports (20, 21, 23, 25, 53, 80, 110, 119, 139, 143, 161).

Fundamentals of TCP/IP [4]

/etc/services
ftp
telnet
host
ping
dig
traceroute
whois

Fundamentals of TCP/IP [4]

TCP/IP Illustrated, Volume 1 The Protocols by W. Richard Stevens Addison Wesley

LPIC 1 Certification Bible Angie Nash and Jason Nash Hungry Minds IP addressing

This objective includes the understanding of IP-addresses, network masks and what they mean (i.e. determine a network and broadcast address for a host based on its subnet mask in "dotted quad" or abbreviated notation or determine the network address, broadcast address and netmask when given an IP-address and number of bits).

IP Address Classes (Classic)

1

\underline{0}0000000.00000000.00000000.00000000 - 0.0.0.0
\underline{0}1111111.11111111.11111111.11111111 - 127.255.255.255

2

\underline{10}000000.00000000.00000000.00000000 - 128.0.0.0
\underline{10}111111.11111111.11111111.11111111 - 191.255.255.255

3

\underline{110}00000.00000000.00000000.00000000 - 192.0.0.0
\underline{110}11111.11111111.11111111.11111111 - 223.255.255.255

IP Address - Loopback Reserved Space 127.0.0.0 — 127.255.255.255 127.0.0.1 localhost

IP Address - Private Networks

There are IP ranges set aside for privite address spaces. These should not be made visible on the internet.

10.0.0.0 — 10.255.255.255
172.16.0.0 — 172.32.255.255
192.168.0.0 — 192.168.255.255

IP Address - Subneting

Network: 192.168.192.0
Subnet:  255.255.255.224

IP Address - Default Route

# route add default gw 192.168.1.1

DoD Layer Model

Application ftp, telnet, mail, http protocols

Transport TCP , UDP protocols

Network IP, ICMP, IGMP protocols

Link Ethernet, Token Ring, FDDI

Basic Internet Protocols

IP

ICMP

TCP

UDP

Ports and Port Numbers

$ less /etc/services
ftp               21/tcp
ftp               21/udp           fsp fspd
ssh               22/tcp           # SSH Remote Login Protocol
ssh               22/udp           # SSH Remote Login Protocol
telnet           23/tcp
telnet           23/udp
# 24 - private mail system
smtp              25/tcp           mail
smtp              25/udp           mail
time              37/tcp           timserver

Ports and Port Numbers

FTP 20, 21 Telnet 23 SSH 22 smtp 25 DNS 53 http 80 pop3 110 nntp 119 netbios 137, 138, 139 imap2 143 snmp 161

Port Number vRanges Original reserved ports (till 1992) (256-1023 UNIX)

Well Known or Famous Port Numbers - Reserved Unprivileged The End

License Of This Document